\begin{thebibliography}{1}

\bibitem{bazres}
VV~Bazhanov and N~Yu Reshetikhin.
\newblock Remarks on the quantum dilogarithm.
\newblock {\em Journal of Physics A: Mathematical and General}, 28(8):2217,
  1995.

\bibitem{dimgar}
Tudor Dimofte and Stavros Garoufalidis.
\newblock Quantum modularity and complex {C}hern-{S}imons theory.
\newblock {\em Communications in Number Theory and Physics}, 12(1):1--52, 2018.

\bibitem{kashaev}
Rinat~M Kashaev.
\newblock A link invariant from quantum dilogarithm.
\newblock {\em Modern Physics Letters A}, 10(19):1409--1418, 1995.

\bibitem{kopp}
{G}ene~{S} {K}opp.
\newblock The {S}hintani--{F}addeev modular cocycle: {S}tark units from $ q
  $-{P}ochhammer ratios.
\newblock {\em 2411.06763v3}, 2025.

\bibitem{manin}
Yu~I Manin.
\newblock Real multiplication and noncommutative geometry (ein {A}lterstraum).
\newblock In {\em The Legacy of Niels Henrik Abel}, pages 685--727. Springer,
  2004.

\bibitem{shintani}
Takuro Shintani.
\newblock On a {Kronecker} limit formula for real quadratic fields.
\newblock {\em J. Fac. Sci., Univ. Tokyo, Sect. I A}, 24:167--199, 1977,
  \href{https://irma.math.unistra.fr/~yalkinog/Shintani.pdf}{https://irma.math.unistra.fr/{$\sim$}yalkinog/Shintani.pdf}.

\bibitem{simon}
Christopher-Lloyd Simon.
\newblock Linking numbers of modular knots.
\newblock {\em arXiv:2211.05957v4}, 2024.

\bibitem{yalkinoglu}
Bora Yalkinoglu.
\newblock A note on {S}hintani's invariants.
\newblock {\em 2408.07309v2}, 2025.

\bibitem{yamamoto}
Shuji Yamamoto.
\newblock Factorization of {S}hintani's ray class invariant for totally real
  fields.
\newblock {\em RIMS Kokyuroku Bessatsu}, 19:249--254, 2010.

\end{thebibliography}

\documentclass[11pt, a4paper]{amsart}
\usepackage{amsmath}
\usepackage{amssymb}
\usepackage{amsthm}
\usepackage[figuresright]{rotating}
\usepackage[all]{xy}
\usepackage{pdfsync}
\usepackage[toc,page]{appendix}
\usepackage{hyperref}
\usepackage{mathtools}
\usepackage{lipsum}
\usepackage[T1]{fontenc}        
\usepackage[utf8]{inputenc}
\usepackage[toc,page]{appendix}
\usepackage{mathtools}
\usepackage{fullpage}

\newcommand{\mat}[1]{\begin{bmatrix}#1\end{bmatrix}}
\newcommand{\ep}{\varepsilon}

\newcommand{\genlegendre}[4]{%
  \genfrac{(}{)}{}{#1}{#3}{#4}%
  \if\relax\detokenize{#2}\relax\else_{\!#2}\fi
}

\newcommand{\eq}[1]{\begin{equation}#1\end{equation}}

\theoremstyle{plain}
\newtheorem{theorem}{Theorem}[section]

\newtheorem{question}{Question}[section]

\newtheorem{remark}{Remark}[section]

\numberwithin{equation}{section}

\begin{document}
\author{Bora Yalkinoglu} 
\address{CNRS and IRMA, Strasbourg}
\email{yalkinoglu@math.unistra.fr}
\date{\today}
\title{Shintani's invariant via cyclic quantum dilogarithm}
\maketitle

\begin{abstract}
We formulate Shintani’s invariant in terms of the cyclic quantum dilogarithm. Building on earlier results that expressed Shintani’s invariant using the $q$-Pochhammer symbol, we show how the cyclic quantum dilogarithm naturally arises in this context, providing new perspectives on the arithmetic significance of Shintani’s construction. 
\end{abstract}

\section{Introduction}
\noindent

\noindent In his seminal paper \cite{shintani}, Shintani introduced certain invariants $X(\mathfrak f)$ in terms of the double sine function and conjectured that these invariants provide abelian extensions for real quadratic number fields, thereby offering a conjectural solution to Hilbert’s 12th problem for these fields. Unfortunately, to this day, Shintani’s beautiful conjecture remains unproven. \\ \\
\noindent The aim of this announcement is to show that Shintani's invariant $X(\mathfrak f)$ can be expressed in terms of the cyclic quantum dilogarithm. In earlier work \cite{yamamoto, yalkinoglu, kopp}, it was established - at increasing levels of generality - that Shintani's invariants admit a description via the $q$-Pochhammer symbol. This was first observed in an example of Yamamoto \cite{yamamoto}, and later proven in full generality by Kopp \cite{kopp}. The connection to the present work comes from the fact that the cyclic quantum dilogarithm appears in the asymptotic expansion of the $q$-Pochhammer symbol at roots of unity (see \cite{bazres}). \\ \\
We begin by recalling the specialized framework of \cite{yalkinoglu}, where the description of $X(\mathfrak f)$ via the $q$-Pochhammer symbol takes a particularly transparent form. We then state our new results, Theorems \ref{newmain} and \ref{newcor}, which express $X(\mathfrak f)$ in terms of the cyclic quantum dilogarithm. In particular, this yields an approximation of $X(\mathfrak f)$ via Kummer extensions of cyclotomic fields. \\ \\
A complete account, including full proofs, will appear in a forthcoming paper.
\section{Shintani's invariant via $q$-Pochhammer symbol}
\subsection{Background}
\noindent Let us quickly recall the framework of \cite{yalkinoglu}: Let $K=\mathbb Q(\sqrt{d})$ be a real quadratic field, denote by $\mathcal O_K$ its ring of integers, by $\mathcal O_K^\times$ its unit group, by $\mathcal O _{K,+}^\times$ its group of totally positive units, by $I_K$ the monoid of non-zero integral ideals of $\mathcal O_K$ and by $\mathrm{Cl}_K(\mathfrak f)$ the (strict) ray class group of conductor $\mathfrak f \in I_K$. \\
We fix a totally positive unit $\varepsilon = \frac{a+b \sqrt{d}}{2}$, with $a,b \in \mathbb N$, which generates $\mathcal O_{K,+}^\times$, and write $\ep' = \tfrac{a-b\sqrt{d}}{2}$ for its conjugate. \\
For each $\mathfrak f \in I_K$, define $g(\mathfrak f) \in \mathbb N$ to be the smallest positive integer such that \eq{\langle \ep^{g(\mathfrak f)}\rangle = (\mathcal O_{K,+}^\times \cap (1+\mathfrak f)).} 
We assume throughout that the minus continued fraction expansion of $\ep$ has length one, i.e., $\ep = [[a]]$.\\ 
If $\mathfrak f = (u+v\sqrt{d}) \in I_K$ (with $u,v \in \mathbb Z$) is principal, Shintani's cone decomposition theorem yields a pair $(x,y) = (x_\mathfrak f,y_\mathfrak f) \in \mathbb Q^2$ defined by \eq{x = [-\tfrac{2 v}{b \, \mathcal N(\mathfrak f)}]_1  \text{ and } y = [\tfrac{bu+av}{b \, \mathcal N(\mathfrak f)}],} where $[ \cdot ]:\mathbb R \to \mathbb R / \mathbb Z$ denotes the fractional part and $[\cdot]_1$ agrees with $[\cdot]$ except that $[0]_1 = 1$. \\ \\
For such principal ideals $\mathfrak f = (u+v\sqrt{d})\in I_K$, we define Shintani's invariant\footnote{Attached to the unit element $1_\mathfrak f \in \mathrm{Cl}_K(\mathfrak f)$.}: \eq{X(\mathfrak f) = X_1(\mathfrak f)X_2(\mathfrak f),} by \eq{ X_1(\mathfrak f) = \prod_{l=1}^{g(\mathfrak f)} \mathcal S(\ep,x_l \ep + y_l) \text{ and } X_2(\mathfrak f) = \prod_{l=1}^{g(\mathfrak f)} \mathcal S(\ep',x_l \ep' + y_l),} where $\mathcal S(\omega,z)$ is the double sine function and the decomposition datum \eq{\{(x_l,y_l)\}_{l=1,..,g(\mathfrak f)} \in {\mathbb Q} ^{ 2 g(\mathfrak f)}} is the (normalized) orbit of $U = \mat{a & -1 \\ 1 &0}$ acting (by matrix multiplication) on $(x,y)$, see \cite{yalkinoglu}. 
 \\
Defining  \eq{\tau_n = U^{\tfrac n 2} \cdot \tfrac{a+i b \sqrt{d}}{2} = \tfrac{T_{n+1}(a)+i b \sqrt{d}}{T_n(a)} \in \mathbb H, \ \ n \in \mathbb Z,} where $T_n(x)$ are the Chebyshev polynomials of the first kind\footnote{Characterized by $T_n(x+x^{-1}) = x^n+x^{-n}$.}, we obtain the discretized modular geodesic $\{\tau_n\}_{n\in \mathbb Z}
\subset \mathbb H$ connecting \eq{\ep = \lim_{n\to\infty} \tau_n \text{ and }\ep' =\lim_{n\to \infty}\tau_{-n}.}
\subsection{Shintani's invariant via $q$-Pochhammer symbol}
The main result of \cite{yalkinoglu} is the following
\begin{theorem}
\label{oldmain}
Assume $\ep = [[a]]$. Let $\mathfrak f = (u+v\sqrt{d}) \in I_K$ be principal, with associated pair $(x,y) \in \mathbb Q^2$ and $g = g(\mathfrak f)$. Then \eq{\label{expression0}X_1(\mathfrak f) = \lim_{n\to \infty}\left\vert \frac{(x,y;\tau_{n-g})_\infty}{(x,y;\tau_{n+g})_\infty} \right\vert \text{, } X_2(\mathfrak f) = \lim_{n\to \infty}\left\vert \frac{(x,y;\tau_{-n-g})_\infty}{(x,y;\tau_{-n+g})_\infty} \right\vert,}
where the $q$-Pochhammer symbol is defined by \eq{(x,y;\tau)_\infty = \prod_{k\geq 0}(1-e^{2\pi i (k \tau + x \tau + y)}).}
\end{theorem}
\noindent Thus, Shintani's invariant $X(\mathfrak f)$ can be approximated along the discrete modular geodesic $\{\tau_n\}_{n\in \mathbb Z}$ by the $q$-Pochhammer symbol. 
\section{Shintani's invariant via cyclic quantum dilogarithm}
\noindent Our new result provides an approximation of Shintani's invariants $X(\mathfrak f)$ along roots of unity using the cyclic quantum dilogarithm \eq{D_{\tfrac m n }(x,y) = \prod_{k=1}^{n-1} (1-e^{2 \pi i (k \tfrac{m}{n} + x \tfrac{m}{n} + y)})^{\tfrac k n },} for $m,n\in \mathbb Z$, $n > 1$ and $(m,n)=1$, which appears in the asymptotic expansion of the $q$-Pochhammer symbol at roots of unity (see, e.g., Proposition 3.2 \cite{bazres}). \\ \\
\noindent Building on earlier work (e.g., Proposition 4.34 \cite{kopp} and Theorem \ref{oldmain} \cite{yalkinoglu}) and exploiting the symmetries of the cyclic (quantum) dilogarithm (cf., \cite{dimgar}), we deduce 
\begin{theorem}
\label{newmain}
Under the assumptions of Theorem \ref{oldmain}, let $\mathfrak t_n = \frac{T_{n-1}(a)}{T_n(a)}$ for $n \in \mathbb N$. Then \eq{X_1(\mathfrak f) = \lim_{n\to\infty}\left\vert \frac{\mathrm{D}_{\mathfrak t_n}(y,x )}{\mathrm{D}_{\mathfrak t_{n+g}}(y,x)}\right\vert  \text{, } X_2(\mathfrak f) = \lim_{n\to\infty}\left\vert \frac{\mathrm{D}_{\mathfrak t_n}(x,y )}{\mathrm{D}_{\mathfrak t_{n+g}}(x,y)}\right\vert .}
\end{theorem}
\noindent Hence, Shintani's invariant is approximated by Kummer extensions of cyclotomic fields. 

\noindent Further, we have 
\begin{theorem}
\label{newcor}
Under the assumptions of Theorem \ref{newmain}, if $\mathfrak f = (u) \in I_K$, we have \eq{X_1(\mathfrak f) = X_2(\mathfrak f) = \lim_{n\to\infty}\left\vert \frac{\mathrm{D}_{\mathfrak t_n}(\tfrac 1 u)}{\mathrm{D}_{\mathfrak t_{n+g}}(\tfrac 1 u)}\right\vert.}
\end{theorem}
\begin{remark}
Numerical computations are in excellent agreement with our theorems.
\end{remark}

\section{Further perspectives}
\noindent The cyclic quantum dilogarithm and modular geodesics (lifting to modular knots) appear prominently in knot theory (see, e.g., \cite{kashaev,simon}). 
\begin{question}
Does Shintani’s invariant admit a natural interpretation in terms of knot theory?
\end{question}
\noindent In light of the quantum five-term relation of the cyclic quantum dilogarithm (and the corresponding quantum five-term relation of the $q$-Pochhammer symbol on the noncommutative torus) explained in \cite{bazres}, we may also ask:
\begin{question}
How are the new formulations of Shintani’s invariant - via the $q$-Pochhammer symbol and the cyclic quantum dilogarithm - connected to Manin’s program on real multiplication \cite{manin}? 
\end{question}

\bibliographystyle{plain}
\bibliography{cyclicshintani}

\end{document}